OpenStudy (anonymous):

f(x)=3x+1 and f-1=x-1/3 then the ordered pair of f(3) help me :(

OpenStudy (anonymous):

\[f(x)=3x+1\] \[f^{-1}(x)=\frac{x-1}{3}\]

OpenStudy (anonymous):

f(3)=3*3+1=10 and f-1(3)=3-1/3=2/3

OpenStudy (anonymous):

\(f(3)=3\times 3+1=9+1=10\) ordered pair is \((3,10)\)

OpenStudy (anonymous):

corresponding ordered pair for \(f^{-1}\) is \((10,3)\)